OpenStudy (anonymous):

Water is being pumped into a tank at a rate P(t), where P is measured in gallons per minute. There is a hole near the bottom of the tank causing the water to leak out at a rate of L(t), where L is also measured in gallons per minute. If the tank initially contained 20 gallons of water, which expression would find the volume of water after 10 minutes?

OpenStudy (linn99123):

Let V(t) be the volume of water at time t, with V(0) = 20. For t >= 0, and assuming that the tank does not empty within the interval t in [0, 10], then we have: V(t) = 20 + int_{x=0}^{x=t}[P(x) - L(x)]dx In general, if it is possible for the tank to empty (which depends on the size of the P(x) and L(x) functions), we would have: V(t) = 20 + int_{x=0}^{x=t}[(P(x) - L(x))*F(x) + max[P(x) - L(x),0]E(x)] dx where: F(x) = 1 if V(x)>0 F(x) = 0 if V(x) =0. E(x) = 1-F(x). Thus, F(x)=1 if and only if the tank is non-empty at time x (and 0 else), and E(x)=1 if and only if the tank is empty at time x (and 0 else). Intuitively, the integral equation above says that V(x) increases with rate P(x) - L(x) at a time x such that V(x)>0, and increases at a rate max[P(x)-L(x),0] if the tank is empty at time x.